# Generative AI in Sociological Research:

### v1

@@ B.2 Weighting
Appendix C Descriptive Bar Charts
@@
Appendix D Regression Model
@@
Trust
Future
Improvement
Net
Positive
@@
Familiarity 0.0440.044* $0.020$ 0.000-0.000 $0.027$ $0.007$ $0.029$
Use $0.138$ $0.727$ 0.659-0.659 $0.534$ $0.455$ $0.652$
Familiarity $\times$ Use $0.035$ $0.064$ 0.0980.098* $0.045$ $0.074$ $0.056$
Not Cis-Man 0.015-0.015 $0.125$ 0.144-0.144 $0.115$ 0.113-0.113 $0.150$
Intercept 1.5661.566*** $0.213$ 4.1004.100*** $0.292$ 2.3252.325*** $0.300$
@@
  • Note: Coefficient estimates are unstandardized. Standard errors are design-adjusted [lumley2017fitting]. $n=$ 411, 407, and 409, respectively, after listwise deletion. Reference category for ”Use” is “Not at least weekly.” Reference category for “Not Cis-Man” is “Cis-Man.” *$p<.05$, **$p<.01$, ***$p<.001$ (two-tailed)
Appendix E Alternative Model Specifications
@@
Trust
Future
Improvement
Net
Positive
@@
Familiarity 0.0960.096* $0.048$ 0.006-0.006 $0.062$ 0.000-0.000 $0.052$
Use $0.787$ $1.830$ 2.260-2.260 $1.420$ $0.461$ $1.190$
Familiarity $\times$ Use $0.033$ $0.157$ 0.2950.295* $0.123$ $0.152$ $0.106$
Not Cis-Man 0.066-0.066 $0.274$ 0.345-0.345 $0.285$ 0.241-0.241 $0.271$
@@
$Y\leq$ Strongly Disagree 0.039-0.039 $0.529$ 5.490-5.490*** $1.140$ 1.180-1.180* $0.564$
$Y\leq$ Somewhat Disagree 2.1102.110*** $0.544$ 3.120-3.120*** $0.719$ $0.191$ $0.541$
$Y\leq$ Neither Agree nor Disagree 3.9603.960*** $0.576$ 1.310-1.310 $0.685$ 1.7301.730*** $0.513$
$Y\leq$ Somewhat Agree 6.1206.120*** $0.849$ $0.405$ $0.677$ 3.3903.390*** $0.580$
@@
$\chi^{2}_{null-model}$ 64.71264.712*** 59.37159.371*** 144.818144.818***
  • Note: Standard errors are design-adjusted [lumley2017fitting]. Estimates are expressed in log odds. $n=$ 411, 407, and 409, respectively, after listwise deletion. Reference category for ”Use” is “Not at least weekly.” Reference category for “Not Cis-Man” is “Cis-Man.” *$p<.05$, **$p<.01$, ***$p<.001$ (two-tailed)

### v2

@@ C.2 Weighting
Appendix D Descriptive Bar Charts
@@
Appendix E Regression Model
@@
Trust Future Improvement Net Positive
@@
Familiarity 0.044 0.020 -0.000 0.027 0.007 0.029
Use 0.138 0.727 -0.659 0.534 0.455 0.652
Familiarity $\times$ Use 0.035 0.064 0.098 0.045 0.074 0.056
Not Cis-Man -0.015 0.125 -0.144 0.115 -0.113 0.150
Intercept 1.566∗∗∗ 0.213 4.100∗∗∗ 0.292 2.325∗∗∗ 0.300
@@
  • Note: Coefficient estimates are unstandardized. Standard errors are design-adjusted [lumley2017fitting]. $n=$ 411, 407, and 409, respectively, after listwise deletion. Reference category for “Use” is “Not at least weekly.” Reference category for “Not Cis-Man” is “Cis-Man.” $^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$ (two-tailed).
Appendix F Alternative Model Specifications
@@
Trust Future Improvement Net Positive
@@
Familiarity 0.096 0.048 -0.006 0.062 -0.000 0.052
Use 0.787 1.830 -2.260 1.420 0.461 1.190
Familiarity $\times$ Use 0.033 0.157 0.295 0.123 0.152 0.106
Not Cis-Man -0.066 0.274 -0.345 0.285 -0.241 0.271
@@
$Y\leq$ Strongly Disagree -0.039 0.529 -5.490∗∗∗ 1.140 -1.180 0.564
$Y\leq$ Somewhat Disagree 2.110∗∗∗ 0.544 -3.120∗∗∗ 0.719 0.191 0.541
$Y\leq$ Neither Agree nor Disagree 3.960∗∗∗ 0.576 -1.310 0.685 1.730∗∗∗ 0.513
$Y\leq$ Somewhat Agree 6.120∗∗∗ 0.849 0.405 0.677 3.390∗∗∗ 0.580
@@
$\chi^{2}_{null-model}$ 64.712∗∗∗ 59.371∗∗∗ 144.818∗∗∗
  • Note: Standard errors are design-adjusted [lumley2017fitting]. Estimates are expressed in log odds. $n=$ 411, 407, and 409, respectively, after listwise deletion. Reference category for “Use” is “Not at least weekly.” Reference category for “Not Cis-Man” is “Cis-Man.” $^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$ (two-tailed).